# A central limit theorem for the stochastic cable equation

### v1

@@ 6 On the limit in Assumption 1
$\mathcal{I}_{k}(t,x):=$ $\sigma_{1}^{k-1}\int_{0}^{t}\int_{0}^{L}\dots\int_{0}^{r_{k-1}}\int_{0}^{L}G_{t-r_{1}}(x,z_{1})\dots G_{r_{k-1}-r_{k}}(z_{k-1},z_{k})$
×$\qquad\qquad\times(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})\ W(\mathrm{d}r_{k},\mathrm{d}z_{k})\dots W(\mathrm{d}r_{1},\mathrm{d}z_{1})$
@@
×$\qquad\qquad\times(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})^{2}\ \mathrm{d}r_{k}\mathrm{d}z_{k}\dots\mathrm{d}r_{1}\mathrm{d}z_{1}$
@@
×$\quad\times\int_{0}^{t}\int_{\mathbb{R}}\dots\int_{0}^{r_{k-1}}\int_{\mathbb{R}}p_{\beta(t-r_{1})}^{2}(x-z_{1})\dots p_{\beta(r_{k-1}-r_{k})}^{2}(z_{k-1}-z_{k})\ \mathrm{d}r_{k}\mathrm{d}z_{k}\dots\mathrm{d}r_{1}\mathrm{d}z_{1}$
@@
×$\quad\times\int_{0<r_{k}<\dots<r_{1}<1}\sqrt{\frac{1}{(1-r_{1})\times\dots\times(r_{k-1}-r_{k})}}\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
@@
×$\quad\times\frac{1}{L}\int_{0}^{L}(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})^{2}\ \mathrm{d}z_{k}\mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
@@
×$\quad\times\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})^{2}\ \mathrm{d}z_{k}\mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
@@
×$\quad\times(\sigma_{1}^{2}e^{-2\alpha r_{k}}+2\sigma_{1}\sigma_{0}e^{-\alpha r_{k}}+\sigma_{0}^{2})\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
@@
×$\quad\times(\sigma_{1}+e^{\alpha r_{k}}\sigma_{0})^{2}\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
@@
×$\quad\times(\sigma_{1}e^{-\alpha r_{k}}+\sigma_{0})^{2}\ \mathrm{d}r_{k}\mathrm{d}z_{k}\dots\mathrm{d}r_{1}\mathrm{d}z_{1}$
@@
×$\quad\times(\sigma_{1}e^{-\alpha r_{k}}+\sigma_{0})^{2}\ \mathrm{d}r_{1}\dots\mathrm{d}r_{k}$
@@
×$\quad\times(\sigma_{1}+\sigma_{0}e^{\alpha r_{k}})^{2}\ \mathrm{d}r_{1}\dots\mathrm{d}r_{k}$
@@
Therefore, in the case of any boundary conditions, we obtain
@@
×$\quad\times\int_{0<r_{k}<\dots<r_{1}<t}p_{2\beta(t-r_{1})}(0)\dots p_{2\beta(r_{k-1}-r_{k})}(0)\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
@@ Appendix A Properties of Green’s function
$G_{t}(x,y)=e^{-\alpha t}\sum_{n\in\mathbb{Z}}\left(p_{\beta t}(x-y+2nL)+p_{\beta t}(x+y+2nL)\right),$
@@
$G_{t}(x,y)=e^{-\alpha t}\sum_{nin\mathbb{Z}}p_{\beta t}(x-y+nL).$

### v2

@@ 6 On the limit in Assumption 1
$\mathcal{I}_{k}(t,x)$= $\sigma_{1}^{k-1}\int_{0}^{t}\int_{0}^{L}\dots\int_{0}^{r_{k-1}}\int_{0}^{L}G_{t-r_{1}}(x,z_{1})\dots G_{r_{k-1}-r_{k}}(z_{k-1},z_{k})$
$\qquad\qquad\times(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})\ W(\mathrm{d}r_{k},\mathrm{d}z_{k})\dots W(\mathrm{d}r_{1},\mathrm{d}z_{1})$
@@
$\qquad\qquad\times(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})^{2}\ \mathrm{d}r_{k}\mathrm{d}z_{k}\dots\mathrm{d}r_{1}\mathrm{d}z_{1}$
@@
$\quad\times\int_{0}^{t}\int_{\mathbb{R}}\dots\int_{0}^{r_{k-1}}\int_{\mathbb{R}}p_{\beta(t-r_{1})}^{2}(x-z_{1})\dots p_{\beta(r_{k-1}-r_{k})}^{2}(z_{k-1}-z_{k})\ \mathrm{d}r_{k}\mathrm{d}z_{k}\dots\mathrm{d}r_{1}\mathrm{d}z_{1}$
@@
$\quad\times\int_{0<r_{k}<\dots<r_{1}<1}\sqrt{\frac{1}{(1-r_{1})\times\dots\times(r_{k-1}-r_{k})}}\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
@@
$\quad\times\frac{1}{L}\int_{0}^{L}(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})^{2}\ \mathrm{d}z_{k}\mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
@@
$\quad\times\lim_{L\to\infty}\frac{1}{L}\int_{0}^{L}(\sigma_{1}\mathcal{I}_{0}(r_{k},z_{k})+\sigma_{0})^{2}\ \mathrm{d}z_{k}\mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
@@
$\quad\times(\sigma_{1}^{2}e^{-2\alpha r_{k}}+2\sigma_{1}\sigma_{0}e^{-\alpha r_{k}}+\sigma_{0}^{2})\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
@@
$\quad\times(\sigma_{1}+e^{\alpha r_{k}}\sigma_{0})^{2}\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
@@
$\quad\times(\sigma_{1}e^{-\alpha r_{k}}+\sigma_{0})^{2}\ \mathrm{d}r_{k}\mathrm{d}z_{k}\dots\mathrm{d}r_{1}\mathrm{d}z_{1}$
@@
$\quad\times(\sigma_{1}e^{-\alpha r_{k}}+\sigma_{0})^{2}\ \mathrm{d}r_{1}\dots\mathrm{d}r_{k}$
@@
$\quad\times(\sigma_{1}+\sigma_{0}e^{\alpha r_{k}})^{2}\ \mathrm{d}r_{1}\dots\mathrm{d}r_{k}$
@@
Therefore, in the case of Neumann, Dirichlet, or periodic boundary conditions, we obtain
@@
$\quad\times\int_{0<r_{k}<\dots<r_{1}<t}p_{2\beta(t-r_{1})}(0)\dots p_{2\beta(r_{k-1}-r_{k})}(0)\ \mathrm{d}r_{k}\dots\mathrm{d}r_{1}$
@@ Appendix A Properties of Green’s function
$G_{t}(x,y)=e^{-\alpha t}\sum_{n\in\mathbb{Z}}\left(p_{\beta t}(x-y+2nL)+p_{\beta t}(x+y+2nL)\right),$ (35)
@@
$G_{t}(x,y)=e^{-\alpha t}\sum_{n\in\mathbb{Z}}p_{\beta t}(x-y+nL).$ (37)